Use substitution or elimination to solve each system of equations.
2x + 5y= 16
5x - 2y=11

Answer by mananth(16949)

2x + 5y= 16..............1
5x - 2y=11...............2
...
multiply (1) by 2 & (2) by 5
we get
4x+10y=32
25x-10y=55
add them
29x=87
/29
x=3
...
plug value of x in (1)
2x+5y=16
2*3+5y=16
6+5y=16
5y=10
/5
y=2